The average weight of 4 football players is 180 pounds. The weights of three of the players are 165 pounds, 185 pounds, and 215 pounds. What is the weight of the fourth player?
Answer by likaaka(51)

To find the average weight, you add the weights and divide the total by the amount of players.
So if you're given the weights of three players, the weight of the fourth will be called x.
(165 + 185 + 215 + x)/4 = 180
Now solve for x
(565 + x)/4 = 180, multiply both sides by 4
((565 + x)/4) * 4 = 180 * 4
565 + x = 720, subtract both sides by 565
565 + x - 565 = 720 - 565
x = 155 so the fourth player weighs 155lbs.
Now check to see if given that the fourth player weighs 155lbs, will the average of all four players be 180lbs.
165 + 185 + 215 + 155 = 720
720/4 = 180